a mississippi riverboat can travel 75 kilometers downstream in three hours and can make the return trip in five hours. find the speed of the boat in still water.
Answer by longjonsilver(2297)

let speed of boat in still water = x
let speed of water = w
so we know:
with water, spped of boat = x+w
against water, speed of boat = x-w

now speed=distance/time

so, with water:
x+w = 75/3
x+w = 25
against water:
x-w = 75/5
x-w = 15

seeing as how we want x, lets get rid of w by adding the 2 equations together to give 2x = 40 --> x = 20km/hr

this is your answer.

jon.